An officer of a certain agency claims that the mean monthly income of a family that lives in a depressed area in a certain town is Php7.500.00. A group of researchers conducted a survey in that area and found out that the mean monthly income of 25 selected families is Php6,000.00 with a standard deviation of Php150.00. Test the claim that = Php7,500.00 at 0.01 level of significance


1
Expert's answer
2021-05-21T14:02:58-0400

Hypothesized Population Mean "\mu=7500" 

Sample Standard Deviation "s=150"

Sample Size "n=25"

Sample Mean "\bar{X}=6000"

Significance Level "\alpha=0.01" 


The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:\mu=7500"

"H_1:\mu\not=7500"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.


Based on the information provided, the significance level is "\alpha=0.01," and "df=n-1=24" degrees of freedom. The critical value for a two-tailed test is "t_c=2.79694." 

The rejection region for this left-tailed test is "R=\{t:|t|>2.79694\}." 


The t-statistic is computed as follows:


"t=\dfrac{\bar{X}-\mu_0}{s/\sqrt{n}}=\dfrac{6000-7500}{150/\sqrt{25}}=-50"



Since it is observed that "|t|=50>2.79694=t_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than "7500," at the "\alpha=0.01" significance level.


Using the P-value approach: The p-value is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than "7500," at the "\alpha=0.01" significance level.